OpenStudy (anonymous):

Kevin plans to put concrete on a rectangular portion of his driveway. The portion is 10 feet long and three inches high. The price of concrete is $98 per cubic yard. The total cost of the concrete Kevin needs is $72.59. Which of the following is closest to the width of the portion of the driveway on which Kevin plans to put concrete? [1 foot = 12 inches; 1 yard = 3 feet] a.) 3 feet b.) 6 feet c.) 8 feet d.) 10 feet

OpenStudy (welshfella):

Cost = volume of concrete needed in cubic yards * 98 72.59 = V * 98

OpenStudy (welshfella):

V in cubic yards = 10 * 1/4 * x ---------- 3^3 where the width x is in feet

OpenStudy (welshfella):

so from the first equation V = 72.59 / 98 = 0.7407 cu yds

OpenStudy (welshfella):

so 0.7407 = 2.5x ----- 363 solve for x

OpenStudy (welshfella):

* typo in the denominator 0.7407 = 2.5 x ----- 3^3

OpenStudy (welshfella):

can you solve that?